OpenStudy (anonymous):

Verify each identity: (sinx-cosx)/(sinx) + (cosx-sinx)/(cosx)=2-secxcscx

OpenStudy (luigi0210):

This looks fun. Any ideas on how to start?

OpenStudy (anonymous):

common denominator?

OpenStudy (luigi0210):

Yes

OpenStudy (luigi0210):

And what do you get?

OpenStudy (anonymous):

\[\frac{ \cos \theta \sin \theta-\cos^2 \theta +\sin \theta \cos \theta -\sin^2 \theta }{ \sin \theta \cos \theta }\]

OpenStudy (luigi0210):

Correct, now combine like terms, and separate the fraction into parts.

OpenStudy (anonymous):

\[\frac{ 2\cos \theta \sin \theta }{ \cos \theta \sin \theta} + \frac{ -\cos^2 \theta - \sin^2 \theta }{ \cos \theta \sin \theta }\]

OpenStudy (luigi0210):

There ya go, now simplify. Take out the negative and make it: \[\LARGE -(\frac{sin^2\theta+cos^2\theta}{cos\theta sin\theta})\] Do you remember what \[\LARGE sin^2+cos^2x=?\]

OpenStudy (anonymous):

1

OpenStudy (luigi0210):

Right. In the first fraction, the cos and sin cancel leaving 2. The sin+cos turn to 1. Now we separate that: \[\LARGE 2+(\frac{1}{cos\theta}*\frac{1}{sin\theta})\] Finish it off~

OpenStudy (anonymous):

\[2-\sec \theta \csc \theta = 2- \sec \theta \csc \theta\]

OpenStudy (luigi0210):

Whoops, forgot my negative, but yea :)

OpenStudy (anonymous):

Thank you so much!!!!! :DDDDDDDDDDD

OpenStudy (luigi0210):

Anytime~